OpenStudy (anonymous):

Some silly questions but i need to be sure about it (sin^2a)^2 (sin^2a +cos^2a)^2 (2sina + 2 cos a )^2 (2sin square a + 2 cos square a ) whole squared (1/3cos^3(9A)) ^2 (tan3A)^2

OpenStudy (anonymous):

\[(\sin^2 a)^2 = \sin^4a\] doesn't it?